OpenStudy (hellokitty17):

Simplify the following expression: (x + 6y) - (3x - 10y). If the final answer is written in the form Ax + By, what is the value of A?

Nnesha (nnesha):

just like last one distribute 2nd parentheses by negative one let me know what you get -1(3x-10y) = ?

oregonduck (oregonduck):

6a + 7b

oregonduck (oregonduck):

(x + 6y) - (3x − 10y)=-2x+16y A=-2

Nnesha (nnesha):

stop giving Direct answers 1) you're not helping him/her 2) you r wasting my time

oregonduck (oregonduck):

what ever